# Corrections to Kerr-Newman from Noncommutative Einstein-Maxwell equation

### v1

@@ IV Solving the NC Einstein-Maxwell equation
$h_{\mu\nu}=\left[\begin{matrix}0&h_{tr}(r,\theta)&h_{t\theta}(r,\theta)&0\\ h_{tr}(r,\theta)&0&0&h_{r\phi}(r,\theta)\\ h_{t\theta}(r,\theta)&0&0&h_{\theta\phi}(r,\theta)\\ 0&h_{r\phi}(r,\theta)&h_{\theta\phi}(r,\theta)&0\end{matrix}\right]\;,$ (24)
but simply using (24) as the $a^{1}$ part of the ansatz (19) and plugging it in (16) will give a system of 4 partial differential equations with each equation being hundreds of terms long. So solving the system of equations in general is computationally intractible  [^6] Calculating the system of equations for the ansatz (24) took two hours on a modern PC using the open source Sympy library. unless one specifies the ansatz in much more detail. The way the author has found the solution was by using physical arguments, not by mechanical computation.
@@
which is not surprising since the theory that we are solving is a $U(1)_{\star}$ noncommutative gauge theory. The metric part of the solution can be obtained by considering the scalar field action (6) in the (commutative) KN background, and in the manner of [7], the author searched for an effective metric $\hat{g}$ for which the Klein-Gordon operator equals the NC scalar field action in the fixed KN background
$\frac{\delta S}{\delta\phi}=\square_{\hat{g}}\phi\;.$ (26)
@@
$\begin{split}\hat{h}_{tr}&=-\frac{qJMQ\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{t\theta}&=-\frac{qJM^{3}Qr\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{r\phi}&=\frac{qQ\left(J^{2}+M^{2}r^{2}\right)\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{\theta\phi}&=\frac{qJ^{2}M^{2}Qr\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;.\\ \end{split}$ (28)
Incredibly, the same metric components (28), up to a factor of $-2$
$h_{\mu\nu}=-2\hat{h}_{\mu\nu}\;,$ (29)
@@
together with the Seiberg-Witten expanded potential (25), solve the noncommutative equations (16) and (18). In conclusion, a charged rotating solution of the NC Einstein-Hilbert-Maxwell action (15) is given as
@@
$\begin{gathered}g_{\mu\nu}=\left[\begin{matrix}\frac{\frac{J^{2}\sin^{2}{\left(\theta\right)}}{M^{2}}-\Delta{\left(r\right)}}{\Sigma{\left(r,\theta\right)}}&-2a\hat{h}_{tr}{\left(r,\theta\right)}&-2a\hat{h}_{t\theta}{\left(r,\theta\right)}&-\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\\ -2a\hat{h}_{tr}{\left(r,\theta\right)}&\frac{\Sigma{\left(r,\theta\right)}}{\Delta{\left(r\right)}}&0&-2a\hat{h}_{r\phi}{\left(r,\theta\right)}\\ -2a\hat{h}_{t\theta}{\left(r,\theta\right)}&0&\Sigma{\left(r,\theta\right)}&-2a\hat{h}_{\theta\phi}{\left(r,\theta\right)}\\ -\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}&-2\hat{h}_{r\phi}{\left(r,\theta\right)}&-2a\hat{h}_{\theta\phi}{\left(r,\theta\right)}&\frac{\left(-\frac{J^{2}\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}}{M^{2}}+\left(\frac{J^{2}}{M^{2}}+r^{2}\right)^{2}\right)\sin^{2}{\left(\theta\right)}}{\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\\ A_{\mu}=\left[\begin{matrix}\frac{Qr}{\Sigma{\left(r,\theta\right)}}\\ 0\\ -\frac{JQ^{2}ar^{2}\sin{\left(2\theta\right)}}{2M\Sigma^{2}{\left(r,\theta\right)}}\\ -\frac{JQr\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\end{gathered}$ (30)
@@
and it satisfies the condition that $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors for $g_{\mu\nu}$ and $A_{\mu}$ so our solution (30) is valid for any action of the form (13) which reproduces the usual Einstein-Hilbert-Maxwell action in the commutative limit $a\rightarrow 0$.
@@
It is worth spending a few words on some properties of the noncommutative solution (30). Firstly, its Komar mass and angular momentum are $M$ and $J$ respectively, so they remain unchanged in the noncommutative spacetime. Secondly, we can see that the dimensional parameter $q$ is present in both the metric and potential part of the solution. This dimensional parameter equals the charge of the scalar field in the theory (6), but since our theory (15) does not have a scalar field, we can not decide on a specific value of the new hair $q$. Notice that in the Seibeg-Witten map (11) the presence of the parameter $q$ is necessary because otherwise the map would not be dimensionally consistent. In the limit $J\rightarrow 0$ we obtain the same metric (up to the $-2$ factor) as the effective metric in [7] where it was also noticed that the effective metric solves the Einstein equation  [^7] In the limit $J\rightarrow 0$ the electromagnetic potential is simple enough that it does not introduce any $a^{1}$ terms to the Einstein equation or the Maxwell equation. up to $\mathcal{O}(a^{2})$.
@@ V Final remarks
In this letter we have explored the inclusion of noncommutative geometry via the $\partial_{t}\wedge\partial_{\phi}$ coordinate Drinfeld twist into the standard Einstein-Hilbert-Maxwell action. We have found that there are infinitely many noncommutative actions which reproduce the EHM action in the commutative limit. That is because in the commutative geometry, there are also infinitely many equivalent EHM actions due to the commutativity of the multiplication, e.g.,
$\mathcal{L}_{1}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}=\mathcal{L}_{2}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}$ (31)
@@
or, e.g., any convex combination of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ gives again an equivalent action. Still, we realized that under sufficiently symmetric conditions, all the NC actions share the equations of motion (16) and (18) which correspond to a nonlinear electrodynamics theory. We have solved those equation of motion and found a family of solutions, parametrized by $q$, which correspond to perturbative corrections to the commutative Kerr-Newman solution, with 4 new metric components, $g_{tr},g_{t\theta},g_{r\phi},g_{\phi\theta}$ appearing as well as one new electromagnetic potential component, $A_{\theta}$, appearing in the solutions. This letter is the first time, to author’s awareness, that the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist defornation formalism was applied to rotating black hole physics. It is also the first time that in this formalism a correction to the black hole physics is obtained directly as a solution to the equation of motion of the NC action and not by some indirect methods. This letter serves as the proof of concept that the quantum nature of spacetime can indeed affect the black hole solutions. Finally, [11] established a result that commutative gravitational solutions with (at least) two Killing vectors are also solutions to the noncommutative Einstein equation constructed from a Killing Drinfeld twist. Here we give more depth to this result by proving that even if the vectors that define the twist are Killing for all of the commutative fields, they can still fail to solve the noncommutative Einstein-matter equation in case the matter content contains gauge fields complicated enough that the Seiberg-Witten map can introduce new nonzero source terms. Never the less, the claim from [11] can be slightly weakened to say that in such cases, the left-hand side of the Einstein equation remains undeformed, it is just the sources which can be deformed from the deformation of gauge symmetry.

### v2

@@ IV Solving the NC Einstein-Maxwell equation
$h_{\mu\nu}=\left[\begin{matrix}0&h_{tr}(r,\theta)&h_{t\theta}(r,\theta)&0\\ h_{tr}(r,\theta)&0&0&h_{r\varphi}(r,\theta)\\ h_{t\theta}(r,\theta)&0&0&h_{\theta\varphi}(r,\theta)\\ 0&h_{r\varphi}(r,\theta)&h_{\theta\varphi}(r,\theta)&0\end{matrix}\right]\;,$ (24)
but simply using (24) as the $a^{1}$ part of the ansatz (19) and plugging it in (16) will give a system of 4 partial differential equations with each equation being hundreds of terms long (without even yet considering the Maxwell part). So solving the system of equations in general is computationally intractable  [^6] Calculating the system of equations for the ansatz (24) took two hours on a modern PC using the open source Sympy library. unless one specifies the ansatz in much more detail. The way the author has found the solution was by using physical arguments, not by mechanical computation.
@@
which is not surprising since the theory that we are solving is a $U(1)_{\star}$ noncommutative gauge theory. The metric part of the solution can be obtained by considering the scalar field action (6) in the (commutative) KN background, and in the manner of [7], the author searched for an effective metric $\hat{g}$ for which the Klein-Gordon operator equals the NC scalar field’s equation of motion in the fixed KN background
$\frac{\delta S}{\delta\phi^{\dagger}}=\square_{\hat{g}}\phi\;.$ (26)
@@
$\begin{split}\hat{h}_{tr}&=-\frac{qJMQ\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{t\theta}&=-\frac{qJM^{3}Qr\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{r\varphi}&=\frac{qQ\left(J^{2}+M^{2}r^{2}\right)\left(J^{2}\cos^{2}{\left(\theta\right)}-M^{2}r^{2}\right)\sin^{2}{\left(\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;,\\ \hat{h}_{\theta\varphi}&=\frac{qJ^{2}M^{2}Qr\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}\sin{\left(2\theta\right)}}{2\left(J^{2}\cos^{2}{\left(\theta\right)}+M^{2}r^{2}\right)^{2}}\;.\\ \end{split}$ (28)
The effective metric (26) for the scalar field reorganizes the NC gauge symmetry of the scalar field and electromagnetic background into the metric effects, but without considering the back-reaction of the gravitational field. In the problem we are solving, the SW expanded $U(1)_{\star}$ potential again drives the reaction and back-reaction. With that said, one can try to see if this back-reaction is of the same form as the effective metric. Following that logic, let us use the effective metric (28) as an ansatz for the Einstein-Maxwell equations (16), (18) by allowing the numerical coefficients of $\hat{h}_{\mu\nu}$ to vary. This amounts to setting the $h_{\mu\nu}$ field in the ansatz (19) to
$h_{\mu\nu}=\left[\begin{matrix}0&C_{1}\hat{h}_{tr}(r,\theta)&C_{2}\hat{h}_{t\theta}(r,\theta)&0\\ C_{1}\hat{h}_{tr}(r,\theta)&0&0&C_{3}\hat{h}_{r\varphi}(r,\theta)\\ C_{2}\hat{h}_{t\theta}(r,\theta)&0&0&C_{4}\hat{h}_{\theta\varphi}(r,\theta)\\ 0&C_{3}\hat{h}_{r\varphi}(r,\theta)&C_{4}\hat{h}_{\theta\varphi}(r,\theta)&0\end{matrix}\right]\;.$ (29)
Incredibly, the ansatz turns out to be functionally correct. This can be seen by plugging it into the Einstein-Maxwell problem (16), (18) which in turn uniquely imposes setting all of the unknown coefficients to
$C_{1}=C_{2}=C_{3}=C_{4}=-2\;.$ (30)
@@
The ansatz (29) (with the restriction (30)) together with the Seiberg-Witten expanded potential (25) solves the NC equations of motion. In conclusion, a charged rotating solution (up to $\mathcal{O}(a^{2})$) of the NC Einstein-Hilbert-Maxwell action (15) is given as
@@
$\begin{gathered}g_{\mu\nu}=\left[\begin{matrix}\frac{\frac{J^{2}\sin^{2}{\left(\theta\right)}}{M^{2}}-\Delta{\left(r\right)}}{\Sigma{\left(r,\theta\right)}}&-2a\hat{h}_{tr}{\left(r,\theta\right)}&-2a\hat{h}_{t\theta}{\left(r,\theta\right)}&-\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\\ -2a\hat{h}_{tr}{\left(r,\theta\right)}&\frac{\Sigma{\left(r,\theta\right)}}{\Delta{\left(r\right)}}&0&-2a\hat{h}_{r\varphi}{\left(r,\theta\right)}\\ -2a\hat{h}_{t\theta}{\left(r,\theta\right)}&0&\Sigma{\left(r,\theta\right)}&-2a\hat{h}_{\theta\varphi}{\left(r,\theta\right)}\\ -\frac{J\left(\frac{J^{2}}{M^{2}}+r^{2}-\Delta{\left(r\right)}\right)\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}&-2\hat{h}_{r\varphi}{\left(r,\theta\right)}&-2a\hat{h}_{\theta\varphi}{\left(r,\theta\right)}&\frac{\left(-\frac{J^{2}\Delta{\left(r\right)}\sin^{2}{\left(\theta\right)}}{M^{2}}+\left(\frac{J^{2}}{M^{2}}+r^{2}\right)^{2}\right)\sin^{2}{\left(\theta\right)}}{\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\\ A_{\mu}=\left[\begin{matrix}\frac{Qr}{\Sigma{\left(r,\theta\right)}}\\ 0\\ -a\frac{qJQ^{2}r^{2}\sin{\left(2\theta\right)}}{2M\Sigma^{2}{\left(r,\theta\right)}}\\ -\frac{JQr\sin^{2}{\left(\theta\right)}}{M\Sigma{\left(r,\theta\right)}}\end{matrix}\right]\;,\end{gathered}$ (31)
@@
and it satisfies the condition that $\partial_{t}$ and $\partial_{\varphi}$ are Killing vectors for $g_{\mu\nu}$ and $A_{\mu}$ so our solution (31) is valid for any action of the form (13) which reproduces the usual Einstein-Hilbert-Maxwell action in the limit  [^7] In other words, the solution (31) is valid for any action which is obtained from the Einstein-Hilbert Maxwell action by introducing the NC geometry and NC gauge theory as outlined in earlies Sections. As discussed in (14), a very big space of NC Lagrangians is compatible with this framework. $a\rightarrow 0$.
@@
It is worth spending a few words on some properties of the noncommutative solution (31). Firstly, its Komar mass and angular momentum are $M$ and $J$ respectively, so they remain unchanged in the noncommutative spacetime. Secondly, we can see that the dimensional parameter $q$ is present in both the metric and potential part of the solution. This dimensional parameter equals the charge of the scalar field in the theory (6), but since our theory (15) does not have a scalar field, we can not decide on a specific value of the parameter $q$. Whether the parameter $q$ is something fundamental like a coupling constant, or if it is a new type of black hole hair, from the presented analyses it is not clear  [^8] It is not unreasonable to imagine that two NC Kerr-Newman solutions could model physical reality with the same commutative hairs $M,J,Q$ but with a different $q$. On the other hand, it also seems plausible that the same $q$ applies to all modifications of commutative solutions, which is already the case for the parameter $a$.. Notice that in the Seiberg-Witten map (11) the presence of the parameter $q$ is necessary because otherwise the map would not be dimensionally consistent. One thing is certain, for any value of $q$ we obtain a solution to the NC Einstein Maxwell problem.
Finally, in the limit $J\rightarrow 0$ we obtain the same metric (up to the already mentioned 2-2- 2 factor) as the effective NC Reissner–Nordström metric in [7] where it was also noticed that the effective metric and commutative electromagnetic potential solve the NC Einstein-Maxwell equation  [^9] In the limit $J\rightarrow 0$ the electromagnetic potential is simple enough that it does not introduce any $a^{1}$ terms to the Einstein equation or the Maxwell equation. The NC Einstein-Maxwell problem in that case is the same as the commutative one. up to $\mathcal{O}(a^{2})$.
@@ V Final remarks
In this letter we have explored the inclusion of noncommutative geometry via the $\partial_{t}\wedge\partial_{\varphi}$ coordinate Drinfeld twist into the standard Einstein-Hilbert-Maxwell action. In the discussion around (14), we have concluded that, for every $\theta^{\mu\nu},$ there are infinitely many equally motivated models of minimally NC deformed Einstein-Maxwell theory. This means that, unless one finds a solution which they all share (for a given $\theta^{\mu\nu}$), exploring physical consequences of such deformations would be troublesome. Afterwards, we have seen that for the $\partial_{t}\wedge\partial_{\varphi}$ twist, all of the minimal NC deformations (13) of the Einstein-Maxwell theory share a solution in case there exists a $\partial_{t},\partial_{\varphi}$ Killing symmetric solution to the equations of motion of (15). In case such a solution to the EOM of (15) exists, it will also satisfy the EOM for all actions (13). The author would like to emphasize how big this space of minimally NC deformed Einstein-Maxwell actions is. For example, in the sense of (14), the reader is welcome to expand the commutative Ricci scalar and to count how many inequivalent orderings of primitive objects there can be upon promoting the commutative $\cdot$ product to the $\star$ product. Additionally, a weighted average of such deformations is also a valid NC deformation. Never the less, all such NC deformations (in the $\partial_{t}\wedge\partial_{\varphi}$ twist) of the Einstein-Hilbert action along with any NC deformation of the Maxwell action will share the solution (31) to the equations of motion.
@@
The solution (31) is actually a family of solutions, parametrized by $q$, which correspond to perturbative corrections to the commutative Kerr-Newman solution, with 4 new metric components, $g_{tr},g_{t\theta},g_{r\varphi},g_{\varphi\theta}$ as well as one new electromagnetic potential component, $A_{\theta}$, appearing in the solutions. Whether the parameter $q$ is more realistically described as some noncommutative hair of the NC Kerr-Newman black hole, or rather as a constant in the model, is currently unresolved. This letter is the first time, to author’s awareness, that the $\partial_{t}\wedge\partial_{\varphi}$ Drinfeld twist deformation formalism was applied to rotating black hole physics. It is also the first time that in this formalism a correction to the black hole physics is obtained directly as a solution to the equation of motion of the NC action and not by some indirect methods. This letter serves as the proof of concept that the quantum nature of spacetime can indeed affect the black hole solutions. Finally, [11] (and further applied in [23]) established a result that commutative gravitational solutions with (at least) two Killing vectors are also solutions to the noncommutative Einstein equation constructed from a Killing Drinfeld twist. Here we give more depth to this result by proving that even if the vectors that define the twist are Killing for all of the commutative fields, they can still fail to solve the noncommutative Einstein-matter equation in case the matter content contains gauge fields complicated enough that the Seiberg-Witten map can introduce new nonzero source terms. Never the less, the claim from [11] can be slightly weakened to say that in such cases, the left-hand side of the Einstein equation remains undeformed, it is just the sources which can be deformed from the deformation of gauge symmetry.